OpenStudy (anonymous):

differentiate the given function: f(x)=\[(3x^3+4x)(x-5)(x+1)\]

OpenStudy (anonymous):

12x^3-24*x^2-62x-20

OpenStudy (anonymous):

that is not it.

OpenStudy (anonymous):

15x^4-48x^3-33*x^2-32*x-20

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

yes it is. not really sure how you got that. can you explain it please?

OpenStudy (anonymous):

were using this equation to find primes, is that correct?

OpenStudy (anonymous):

sure (3x^3+4*x)*(x-5)(x+1) multiply (3x^3+4*x)*(x-5) = 3x^4 - 15x^3 + 4x^2 - 20x take the previous (3x^4 - 15x^3 + 4x^2 - 20x) (x+1) 3x^5 - 15x^4 + 4x^3 - 20x^2 - 3x^4 - 15x^3 + 4x^2 - 20x just add them then diff

OpenStudy (anonymous):

@surdawi, that is an incorrect method, USE PRODUCT RULE

OpenStudy (anonymous):

and chain rule if applicable

OpenStudy (anonymous):

hey psi, can i say that f(x)=(3x^3+4x) ?

OpenStudy (anonymous):

and then find the prime of that?

OpenStudy (anonymous):

so f prime would be 9x^2+4?

OpenStudy (anonymous):

nope keep f(x) as it is , i.e. as given in the question to find the derivative you have f'(x)= (6x+4)*(x-5)(x+1)+(3x^2+4x) * (1-0)*(x+1)+(3x^2+4x) *(x-5)*(1+0)

OpenStudy (anonymous):

we used chain and product rule we find derivatives of each function at a time i.e.DER OF (3x^2+4x)=6x+4, multiplied BY OTHER ORIGINAL FUNCTION DER OF (X-5)=1, MULTIPLIED BY OTHER ORIGINAL FUNCTIONS DER OF (X+1)=1, MULTIPLIED BY OTHER ORIGINAL FUNCITONS

OpenStudy (anonymous):

@psi9epsilon, you still will get my answer 15x^4-48x^3-33*x^2-32*x-20

OpenStudy (anonymous):

@surdawi, certainly we will still get the answer, but you need to learn how to solve the problem correctly and efficiently : ), this matters most when you are helping someone